# Retrieval-Enhanced Knowledge Editing for Multi-Hop Question Answering in Language Models

### v1

@@ 4.1.2. Editing Baselines
For comparison, we consider three kinds of model editing methods: (1) Model weight updating methods: Fine-tuning (Zhu et al., 2020) edits the entire model weights by language modeling the edited knowledge. ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b) focus on identifying and updating particular neurons associated with the knowledge that needs editing. (2) Auxiliary models methods: SEARC (Mitchell et al., 2022) trains an extra language model to store updated knowledge, and it switches to the auxiliary model when answering questions relevant to the edited facts. Details about the auxiliary models for different language models are included in Appendix B. (3) In-context learning-based method: Mello (Zhong et al., 2023) edits model outputs with multi-round conversations.
Table 2. Edited accuracy (%) on MQUAKE datasets (MQUAKE-CF and MQUAKE-T).
@@
Language Models Datasets
w/o edit
(orig ans)
w/o edit
(edited ans)
Fine
Tuned
ROME MEMEIT SEARC Mello Ours
Models w/o Instruction Tuning GPT-2 MQUAKE-CF 10.7 0.0 3.8 1.7 2.3 4.0 0.0 62.8
MQUAKE-T 4.7 0.0 5.8 6.4 1.6 2.7 0.0 61.8
GPT-J MQUAKE-CF 18.1 0.0 7.7 7.6 8.1 6.8 15.3 69.3
MQUAKE-T 15.3 3.1 3.1 4.1 10.6 2.8 36.7 63.9
Falcon MQUAKE-CF 33.7 1.0 5.6 1.7 2.3 7.9 10.7 66.8
MQUAKE-T 35.9 8.9 17.2 7.3 1.6 4.5 51.5 61.6
Models w/ Instruction Tuning Vicuna MQUAKE-CF 43.2 1.0 4.8 8.4 7.6 7.9 10.2 67.2
MQUAKE-T 15.3 3.1 23.1 5.0 1.7 4.5 51.7 63.2
Llama2 (chat) MQUAKE-CF 29.0 1.1 5.4 6.3 3.8 7.9 20.7 69.1
MQUAKE-T 26.7 5.1 17.1 8.7 1.7 4.5 49.4 66.2
@@ 4.1.3. Implementation Details
We evaluate our editing method on the MQUAKE-CF and MQUAKE-T datasets (Zhong et al., 2023), comprising 1000 counterfactual editing instances per 2-hop, 3-hop, and 4-hop questions in MQUAKE-CF, and total 1868 editing instances for 2-hop and 3-hop questions in MQUAKE-T. In our setting, we edit all these facts at the same time. Following previous work (Zhong et al., 2023), we leverage cases from the MQUAKE-CF-9k dataset to craft prompt templates for both baselines and our method. An edit is considered successful if the edited LLM correctly predicts the updated answer within the first ten generated tokens. This criterion is reported as edited accuracy in Table 2. More setting details are in Appendix. B.
@@ 4.2. Editing Performance Evaluation
To answer Q1, we assess our model editing method across various language models, compared against different baseline methods. We focus on the edited accuracy of these models in answering multi-hop questions with edited answers. For references, we also report the accuracy of un-unedited models in predicting both original and edited answers. Our key observations are: (1) Our RAE outperforms all others in both datasets across five language models, achieving an average improvement of 40.57% over the second-best method (Mello) when conducting thousands of edits at the same time. This superior performance primarily stems from our novel MI-based retrieval objective and an effective pruning strategy. Our design can also seamlessly integrate an external knowledge graph, which effectively links all edited facts, thereby facilitating the multi-hop editing process. (2) Mello demonstrates good performance on the MQUAKE-T dataset with models larger than 6B. However, it underperforms on the MQUAKE-CF dataset and fails with GPT-2, likely due to GPT-2’s inability to follow Mello’s complex prompts, resulting in no output. (3) Other methods generally show lower performance across all language models, which aligns with the findings from MQUAKE dataset paper (Zhong et al., 2023).
Table 3. Multi-hop facts retrieval accuracy (%) comparison.
MQUAKE-CF (300 cases for each kind of question)
@@
Probability MaxProb (GPT-2) 77.7 50.3 67.3 25.3 65.0 20.0
MaxProb (Llama2) 78.3 55.7 79.7 37.0 69.3 28.7
@@ 4.3.1. Retrieval Baselines
We consider two types of retrieval methods as baselines, namely the embedding-based and the probability-based methods. We introduce their details as follows.
(1) Embedding-based retrieval methods mostly concentrate on text retrieval rather than triplet retrieval. For a fair comparison, we adapt them to fit our task by employing the Contriever (Izacard et al., 2021) to encode all edited facts into embeddings, and then cache these embeddings for similarity search. Different retrieval methods use distinct strategies to modify the original $K$-hops questions as the final query to perform the search via dot product between the embeddings.
@@
  • Question Reform (Yadav et al., 2020; Nie et al., 2020) appends the retrieved entity to the previous query for the next hop retrieval. This design is motivated by simulating a reasoning path where the retrieved fact at each hop is viewed as an intermediate reasoning result.
  • Question Decompose, proposed by Mello (Zhong et al., 2023), decomposes a multi-hop question into multiple sing-hop questions. Specifically, a comprehensive conversational-style prompt is applied to interact with the LLM for collecting each single-hop question. For each single-hop question, the corresponding knowledge is retrieved with the naïve method described in KG Link.
(2) In contrast with the embedding-based methods that estimate the similarities between the query and the facts with the dot products of their embeddings, probability-based methods leverage language models to directly estimate this similarity.
  • Max Prob (Zhang et al., 2022; Sun et al., 2023) retrieves the subgraph $G_{S}\in\mathcal{G}^{*}$ that maximizes the conditional probability $p(G_{S}|q)$, where $q$ is the input multiple question. In our experiments, we extract a $K$-hop subgraph from the edited knowledge graph $\mathcal{G}^{*}$ and estimate the probability with a language model $f_{\phi}$ as shown in Section 3.2.3.
@@ 4.3.2. Implementation Details
We select 300 cases for each of the 2, 3, and 4-hop problems from the MQUAKE-CF dataset. For each type, we report two matching accuracy scores in Table 3. Specifically, within the retrieved facts, if any fact also appears in the fact chain of a multi-hop question, we name it a partial match (PM). And if every retrieved fact matches the fact chain for the multi-hop question, we name it exact match (EM).
@@ 4.3.3. Results
We have the following observations: (1) Our proposed mutual information-based retrieval method demonstrates excellent performance across various LLMs in multi-hop fact extraction. We also find its success with relatively small language models, such as GPT-2, showing strong generalization ability. (2) In contrast, traditional embedding-based methods (KG Link and Question Reform) underperform in this multi-hop fact retrieval challenge. Their limitation mainly lies in failing to comprehend the complex interplay between multiple relations in a question. Thus, it hinders the extraction of target facts from the extensive knowledge base. (3) Mello demonstrates the efficacy of decomposing multi-hop questions into single-hop questions for this multi-hop facts retrieval task. Notably, its EM performance drops significantly with an increasing number of hops, probably because it becomes much more challenging for language models to perform question decomposing. (4) Probability maximization-based methods outperform traditional embedding methods, while there is still a gap compared to ours, probably because the predicted most probable fact may not be the most necessary one for answering a specific question.
@@
Improved 10.3 $\uparrow$ 11.8 $\uparrow$ 9.3 $\uparrow$ 9.7 $\uparrow$ 5.7 $\uparrow$
@@
Improved 10.1$\uparrow$ 11.6$\uparrow$ 6.5 $\uparrow$ 7.7 $\uparrow$ 5.5$\uparrow$
@@
Improved 12.0 $\uparrow$ 8.1 $\uparrow$ 7.7 $\uparrow$ 4.0 $\uparrow$ 4.2 $\uparrow$
@@ 4.4. Ablation Studies on Pruning Strategy
Table 4 reports the edited accuracy of RAE working with or without the pruning strategy, and we draw the following conclusions: (1) The proposed pruning technique significantly enhances the performance of model editing, demonstrated by achieving an average accuracy improvement of 8.3% across various language models. (2) We observe a more profound improvement for smaller language models on complex questions, while this benefit to larger language models is relatively less significant. In particular, the performance improvement of GPT-2 on the 4-hop questions reaches 12.0%, while for Llama-2 it is just 4.2%. This observation suggests that larger models are more robust to redundant information.
@@
Figure 4. Editing performance compared with inference cost over different proprietary models.
@@ 4.5. Editing Performance on Proprietary LLMs
To answer Q4, we apply the proposed RAE to edit proprietary language models, where we can only access the model via APIs, such as ChatGPT (OpenAI, 2023a). In such cases, RAE utilizes a different lightweight language model to perform relevant facts retrieval and pruning as introduced in Section 3.2 and Section 3.3. Then, the obtained facts will be fed into the proprietary models to perform in-context editing. We evaluate our proposed editing method on GPT-babbage-002, GPT-3.5-turbo-0613, GPT-3.5-instruct, and GPT-4-0613 models (OpenAI, 2023b). We use GPT-2 (1.3B) as our retrieval model. We report the edited accuracy and total editing cost of our method for randomly selected 300 cases (MQUAKE-CF) in Figure 4. The editing cost includes the total fees of calling APIs and the cost of running GPT-2 for knowledge retrieval on rented GPUs. We also report the results of Mello (Zhong et al., 2023) for comparison, where only the API fees are counted.
In Figure 4, we first observe that Babbage has 0.0% edited accuracy by using Mello, showing its ineffectiveness in editing the un-instruction tuned proprietary language model. It is expected since Mello relies on the conversational ability of language models to decompose multi-hop questions. In contrast, RAE is effective in editing all these proprietary models with a remarkably lower cost than Mello. In particular, RAE improves Mello for editing GPT-4 with almost 20% edited accuracy by only costing around its 15% budget. This highlights the benefit of utilizing the inherent language modeling ability instead of the instruction following ability to perform knowledge retrieval for multi-hop question answering.
@@ 4.6. Case Study
In Figure 5, we present two cases from the MQUAKE-CF dataset to demonstrate the retrieving process over a knowledge graph and the pruning process of the retrieved facts. For visualizations, the red, black, and dotted lines represent the final, candidate, and discarded paths in the knowledge graph with beam search, reflecting the decision-making process of our retrieval design. In Figure 5(a), a potential path Misery–language–¿English–country–¿U.K. is discarded even though it can lead to the correct answer (U.K. in this case). This is because it does not share the information that is needed to answer the target question, resulting in a low MI score. In Figure 5(b), even though this question is associated with three edited facts, our method can successfully locate all of them, demonstrating robustness over multi-hop questions. In both cases, our pruning strategy successfully truncates the retrieved fact chain to only contain necessary facts relying on the normalized model editing entropy.
@@
Figure 5. Case studies for edited facts retrieval and pruning. The retrieval process involves the beam search, starting from the query entity and navigating through the knowledge graph with two beams. The two primary candidate edges (beams) at each entity hop are boldfaced, and the others are discarded and denoted with dashed lines. We emphasize the final beam search result with red color.
5. Related Work
5.1. Model Editing
Existing model editing methods for single-hop fact editing can be categorized into the following kinds: (1) Meta-learning method: MEND (Mitchell et al., 2021) employs an auxiliary network for gradient transformation during fine-tuning. (2) Local modification methods: MEMIT (Meng et al., 2022b) and ROME (Meng et al., 2022a) edit specific neurons within models with minimal impact on unrelated weights. (3) Extension-based methods: SERAC utilizes a fine-tuned small auxiliary language model to answer questions that require edited facts. Meanwhile, GRACE (Hartvigsen et al., 2022) provides a theoretical basis for concept erasure across model layers. (4) In-context learning methods: IKE (Zheng et al., 2023) edit models through in-context learning, bypassing the need for parameter updates.
In contrast, model editing methods for multi-hop fact editing remain underexplored. One such method, Mello (Zhong et al., 2023), utilizes a comprehensive prompt to decompose the multi-hop question into single-hop questions and then retrieve edited knowledge iteratively. A recent contemporaneous work, DeepEdit (Wang et al., 2024), edits LLMs through decoding with constraints, where a neuro-symbolic method is proposed for better reason coherence. However, these approaches heavily rely on the strong instruction following and reasoning ability of LLMs, which refers to their capacity to understand and follow the input instruction to generate coherent, context-aware responses. Existing methods (Zhong et al., 2023) fall short for models in relatively small sizes, such as GPT-J(6B) or Vicuna(7B), compared to powerful GPT-3.5.
5.2. Multi-hop QA with Knowledge Graphs
Multi-hop Knowledge Graph Question Answering (KGQA) involves locating answers within a KG that are several hops away from the starting question entities. Existing techniques usually first extract a relevant subgraph from the entire KG, then apply multi-hop reasoning (Sun et al., 2018; Zhang et al., 2022). Initial retrieval strategies apply similarity scores like PageRank (Sun et al., 2018) or embedding similarity (Zhang et al., 2022; Jiang et al., 2022b) for subgraph selection, but these methods often fail to grasp the logic of the question, potentially overlooking vital facts. In the reasoning stage, earlier research used graph network architectures for reasoning (Yasunaga et al., 2021; Jiang et al., 2022a). Some recent approaches now employ LLMs for reasoning (Wang et al., 2023b; Sun et al., 2023). Think-on-graph applies the LLM as an interactive agent on KGs, using retrieved knowledge for reasoning (Sun et al., 2023). However, these methods rely on strong reasoning capabilities, while LLMs with less strong reasoning ability tend to underperform with these methods.

### v2

@@ 4.1.2. Editing Baselines
For comparison, we consider three kinds of model editing methods: (1) Model weight updating methods: Fine-tuning (Zhu et al., 2020) edits the model weights by language modeling the edited knowledge. ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b) focus on identifying and updating particular neurons associated with the knowledge that needs editing. (2) Auxiliary models methods: SEARC (Mitchell et al., 2022) trains an extra language model to store updated knowledge, and it switches to the auxiliary model when answering questions relevant to the edited facts. (3) RAG-based methods: Mello (Zhong et al., 2023) and DeepEdit (Wang et al., 2024) represent cutting-edge editing methods for multi-hop questions, employing multi-round conversations to edit model outputs. Additionally, the Subgraph Retriever (SR) (Zhang et al., 2022) introduces an advanced knowledge retrieval approach for multi-hop question-answering tasks. We adapt their retrieval method as a baseline.
Table 2. Edited accuracy (%) on multi-hop question editing datasets.
@@
Language Models Datasets Fine Tune ROME MEMIT SEARC Mello DeepEdit
Subgraph
Retriever
RAE(ours)
GPT-2 (1.5B) M-CF 3.8 1.7 2.3 4.0 0.0 0.0 21.9 62.8
M-T 5.8 6.4 1.6 2.7 0.0 0.0 20.3 61.8
Popular 6.2 4.3 2.9 1.1 0.0 0.0 26.7 47.1
GPT-J (6B) M-CF 7.7 7.6 8.1 6.8 15.3 9.3 36.2 69.3
M-T 3.1 4.1 10.6 2.8 36.7 19.6 51.2 63.9
Popular 6.8 7.5 4.4 1.3 12.8 6.6 45.8 49.6
Falcon (7B) M-CF 5.6 1.7 2.3 7.9 10.7 10.8 40.1 66.8
M-T 17.2 7.3 1.6 4.5 51.5 31.7 56.1 61.6
Popular 2.1 4.0 1.1 3.0 8.1 9.5 43.0 50.0
Vicuna (7B) M-CF 4.8 8.4 7.6 7.9 10.2 11.4 39.4 67.2
M-T 23.1 5.0 1.7 4.5 51.7 40.4 58.6 63.2
Popular 4.0 3.8 2.4 3.0 7.7 8.2 29.5 36.1
Llama2 (chat) (7B) M-CF 5.4 6.3 3.8 7.9 20.7 11.2 45.7 69.1
M-T 17.1 8.7 1.7 4.5 49.4 37.9 63.1 66.2
Popular 5.2 13.8 4.9 3.0 13.5 11.1 41.9 51.4
@@ 4.1.3. Implementation Details
We evaluate our editing method on the MQUAKE-CF and MQUAKE-T datasets from (Zhong et al., 2023) and Popular datasets from  (Cohen et al., 2023). The MQUAKE-CF (M-CF) comprises counterfactual editing instances in 2, 3, and 4-hop questions, totaling 3000 edits. The MQUAKE-T dataset (M-T) features temporal editing examples in 2 and 3-hop questions, with a total of 1868 edits. Additionally, the Popular dataset contains counterfactual editing in 2-hop questions, comprising 274 edits. Following previous work (Zheng et al., 2023; Zhong et al., 2023), we leverage relevant cases from the MQUAKE-CF-9k dataset to craft prompt templates for both baselines and our method. We evaluated our editing method using the multi-hop edited accuracy metrics from (Zhong et al., 2023; Wang et al., 2024). The results, which reflect the accuracy when all edits are applied in one batch, are reported in Table 2.
@@ 4.2. Editing Performance Evaluation
To answer Q1, we assess our model editing method across various language models, compared against different baseline methods. Our key observations from Table 2 are: (1) Our RAE outperforms all others in three datasets across five language models when conducting thousands of edits at the same time. This superior performance primarily stems from our novel MI-based retrieval objective and an effective pruning strategy. Our design can also seamlessly integrate an external knowledge graph, which effectively links all edited facts, thereby facilitating the multi-hop editing process. (2) RAG-based methods generally show better performance than other methods. Specifically, Mello and DeepEdit demonstrate good performance on the M-T dataset with models larger than 6B. However, they underperform on the M-CF and popular datasets and fail with GPT-2, likely due to GPT-2’s inability to follow their complex prompts, resulting in no output. Additionally, the subgraph retriever ranks second in performance. However, its probability-driven retrieval method occasionally fails to fetch relevant facts, leading to reduced effectiveness. (3) Other methods generally show lower performance across all language models, which aligns with the findings from (Zhong et al., 2023).
Table 3. Multi-hop facts retrieval precision (%) comparison.
MQUAKE-CF
@@
Probability SR(GPT-2) 77.7 50.3 67.3 25.3 65.0 20.0
SR(Llama2) 78.3 55.7 79.7 37.0 69.3 28.7
@@ 4.3.1. Retrieval Baselines
We consider three embedding-based and one probability-based methods as baselines.
(1) Embedding-based methods mostly utilize dense retrieval to fetch relevant corpus from vector databases. Here, we adapt them to fit our multi-hop knowledge graph retrieval task. We employ a cutting-edge retriever, the Contriever (Izacard et al., 2021), to encode all edited facts into embeddings, and then cache these embeddings for similarity search. In addition to Mello, we include two representative multi-hop retrieval methods as follows:
@@
  • Question Reform (QR) (Yadav et al., 2020; Nie et al., 2020) appends the retrieved entity to the previous query for the next hop retrieval. This design is motivated by simulating a reasoning path where the retrieved fact at each hop is viewed as an intermediate reasoning result.
(2) Probability-based method (i.e., subgraph retriever (Zhang et al., 2022)) retrieves the subgraph $G_{S}\in\mathcal{G}^{*}$ that maximizes the conditional probability $p(G_{S}|q)$, where $q$ is the input multi-hop question.
@@ 4.3.2. Implementation Details
Following (Schütze et al., 2008), we select 300 cases for each of the 2, 3, and 4-hop questions from the M-CF dataset. For each type, we report the retrieval precision scores in Table 3. Specifically, we use the metric $Precision@K$, which calculates the proportion of relevant facts within the top $K$ results (Schütze et al., 2008). Here, $Precision@K=|\{\text{relevant facts}\}|/{K}\times 100\%$, abbreviated as $P@K$. For this study, a $k$-hop question has $k$ facts as its fact chain. A fact is considered relevant if it is part of the fact chain used to answer a multi-hop question.
@@ 4.3.3. Results
We have the following observations: (1) Our proposed mutual information-based retrieval method demonstrates excellent performance across various LLMs in multi-hop fact extraction. We also find its success with relatively small language models, such as GPT-2, showing strong generalization ability. (2) In contrast, traditional embedding-based methods (KG Link and Question Reform) underperform in this multi-hop fact retrieval challenge. Their limitation mainly lies in failing to comprehend the complex interplay between multiple relations in a question. Thus, it hinders the extraction of target facts from the extensive knowledge base. (3) Mello demonstrates the efficacy of decomposing multi-hop questions into single-hop questions for this multi-hop facts retrieval task. Notably, its performance drops significantly with an increasing number of hops, probably because it becomes much more challenging for language models to perform question decomposing. (4) Probability maximization-based methods outperform traditional embedding methods, while there is still a gap compared to ours, probably because the predicted most probable fact may not be the most necessary one for answering a specific question.
@@
Gain 16.3$\uparrow$ 18.5$\uparrow$ 14.3$\uparrow$ 15.2$\uparrow$ 8.1$\uparrow$
@@
Gain 23.4$\uparrow$ 21.6$\uparrow$ 11.7$\uparrow$ 14.0$\uparrow$ 9.1$\uparrow$
@@
Gain 24.0$\uparrow$ 13.8$\uparrow$ 13.9$\uparrow$ 6.5$\uparrow$ 6.8$\uparrow$
@@ 4.4. Ablation Studies on Pruning Strategy
Table 4 reports the edited accuracy of RAE working with or without the pruning strategy, and we draw the following conclusions: (1) The proposed pruning technique significantly enhances the performance of model editing, demonstrated by achieving an average accuracy improvement of 14.5% across various language models. (2) We observe a more profound improvement for smaller language models on complex questions, while this benefit to larger language models is relatively less significant. In particular, the performance improvement of GPT-2 on the 4-hop questions reaches 24.0%, while for Llama-2 it is just 6.8%. This observation suggests that larger models are more robust to redundant information.
@@
Figure 4. Editing performance and inference cost over different proprietary models.
@@ 4.5. Editing Performance on Proprietary LLMs
To answer Q4, we apply the proposed RAE to edit proprietary language models, where we can only access the model via APIs, such as ChatGPT (OpenAI, 2023a). In such cases, RAE utilizes a different lightweight language model to perform relevant facts retrieval and pruning as introduced in Section 3.2 and Section 3.3. Then, the obtained facts will be fed into the proprietary models to perform in-context editing. We evaluate our proposed editing method on GPT-babbage-002, GPT-3.5-turbo-0613, GPT-3.5-instruct, and GPT-4-0613 models (OpenAI, 2023b). We use GPT-2 (1.5B) as our retrieval model. We report the edited accuracy and total editing cost of our method for 300 randomly selected cases (MQUAKE-CF) in Figure 4. The editing cost includes the total fees of calling APIs and the cost of running GPT-2 for knowledge retrieval on rented GPUs. We also report the results of Mello (Zhong et al., 2023) for comparison, where only API fees are counted.
In Figure 4, we first observe that Babbage has 0.0% edited accuracy by using Mello, showing its ineffectiveness in editing the un-instruction tuned proprietary language model. This is expected since Mello relies on the conversational ability of language models to decompose multi-hop questions. In contrast, RAE is effective in editing all these proprietary models with a remarkably lower cost than Mello. In particular, RAE improves Mello for editing GPT-4 with almost 20% edited accuracy by only costing around its 15% budget. This highlights the benefit of utilizing the inherent language modeling ability instead of the instruction following ability to perform knowledge retrieval for multi-hop question answering.
4.6. Performance with Different Batch Sizes
In this section, we evaluate editing performance with different editing batch sizes. Specifically, we set the editing instances in sizes of 1, 10, 100, and 1000 cases for 2-hop questions. We choose Mello (Zhong et al., 2023) as our baseline. The result is presented in Figure 5. We can observe that, in both the Vicuna and Llama2 models, RAE’s accuracy remains stable across different editing instances, whereas Mello’s accuracy significantly declines with increasing instances.
Figure 5. Edited accuracy with different edit batch sizes.
@@ 4.7. Case Study
In Figure 6, we present two cases from the M-CF dataset to demonstrate the retrieval process on a knowledge graph and the pruning process of the retrieved facts. For visualizations, the red, black, and dotted lines represent the final, candidate, and discarded paths in the knowledge graph with beam search, reflecting the decision-making process of our retrieval design. In Figure 6(a), a potential path Misery–language–¿English–country–¿U.K. is discarded even though it can lead to the correct answer (U.K. in this case). This is because it does not share the information that is needed to answer the target question, resulting in a low MI score. In Figure 6(b), even though this question is associated with three edited facts, our method can successfully locate all of them, demonstrating robustness over multi-hop questions. In both cases, our pruning strategy successfully truncates the retrieved fact chain to only contain necessary facts relying on the normalized model editing entropy.
@@
Figure 6. Case studies for edited facts retrieval and pruning. The retrieval process involves the beam search, starting from the query entity and navigating through the knowledge graph with two beams. At each entity hop, the two primary candidate edges are highlighted in bold, while others are discarded and marked with dashed lines. The beam search result with the highest MI score is emphasized in red.
5. Related Work: Model Editing
Existing editing methods can be categorized into the following kinds: First, methods that alter model parameters, including fine-tuning (Zhu et al., 2020), locate-then-edit (Meng et al., 2022a, b), and meta-learning (Mitchell et al., 2021; Zhang et al., 2024b), are prone to catastrophic forgetting issue (Dai et al., 2021; Zhai et al., 2023; Cohen et al., 2023), where the previously encoded knowledge could be lost after editing. They also struggle with using updated language for knowledge reasoning, which can lead to severe hallucinations. (Gekhman et al., 2024; Zhong et al., 2023; Chen et al., 2024). While these methods perform well in single-hop editing, they are less effective in multi-hop scenarios. (Zhong et al., 2023). Second, methods that depend on training auxiliary models also fall short in these scenarios. The auxiliary models are usually smaller language models, which lack the necessary reasoning capability to infer correct answers (Mitchell et al., 2022). In contrast, a third category of methods, based on Retrieval-Augmented Generation (RAG), modifies model outputs in a more effective manner (Zhong et al., 2023; Yin et al., 2024; Wang et al., 2024; Gu et al., 2024b). These methods integrate updated knowledge directly into the model input and edit LLMs through an editing prompt. (Zheng et al., 2023; Cohen et al., 2023). RAG-based approaches offer notable benefits as they are resistant to catastrophic forgetting, and allow for on-the-fly editing. (Wang et al., 2023; Han et al., 2023). For multi-hop editing, existing RAG-based methods perform well on powerful proprietary LLMs like GPT-3.5, but show significant degradation on less robust LLMs such as Llama2-7b, particularly as the number of editing instances increases. In contrast, our RAE can maintain good editing performance even with relatively small LLMs and large editing batch sizes.